# First Statistical Detection of Cool Gas Outflows with JWST Towards Cosmic Dawn

### v1

@@ 4.3 Scaling Relations of Outflow Properties
4.4 Synergy of Stacking and Individual Analysis
@@
A key strength of our work is the synergistic combination of a statistical stacking analysis with a targeted study of individual detections. As Figure 8 in Appendix B illustrates, the galaxies with the most prominent, high-S/N absorption features are typically massive ($M_{*}>10^{10}~\rm{M_{\odot}}$). These individual detections serve as unambiguous proof-of-concept, confirming the physical reality of the outflow phenomenon and revealing the rich diversity of absorption profile morphologies. This diversity includes not only powerful outflows but also a few intriguing cases of redshifted absorption as potential signatures of gas accretion (inflow), which we defer to a future, more detailed analysis (see Appendix B).
However, these massive objects represent only a fraction of our full sample. The vast majority of galaxies, particularly at lower stellar masses ($M_{*}<10^{9.5}~\rm{M_{\odot}}$), are numerous but have spectra with insufficient S/N for individual analysis. Thanks to the S/N-weighted stacking technique, we can unlock the statistical power of this dominant population, recovering the faint, average outflow signatures that are otherwise invisible. By bridging these two regimes, our approach provides the most comprehensive view possible, leveraging every spectrum to its full potential and robustly characterizing the properties of galactic outflows across the full dynamic range of our sample.
5 Summary
@@
We present the first statistical study of cool gas outflows in galaxies spanning a wide cosmic timeline from $z\approx 1$ to $z\approx 10$. Leveraging thousands of public JWST/NIRSpec spectra, we employ a S/N-weighted spectral stacking technique, complemented by an analysis of individual high-S/N detections, to robustly characterize the faint absorption signatures of the Mg II doublet. Our sample is shown to be broadly representative of the underlying galaxy population. Our main findings, derived from this powerful combined approach, are summarized as follows:
@@
    We robustly detect blueshifted Mg II absorption, indicative of cool galactic outflows, in all redshift and stellar mass bins across our entire sample range ($1<z<10$). The typical outflow velocities are substantial ($\sim 350~{\rm km~s^{-1}}$), demonstrating that powerful outflows are a persistent feature of galaxy evolution. This work presents the first statistical detection of cool gas outflows in galaxies at $z>3$.
@@
    The properties of cool gas outflows are linked to stellar mass of the host galaxy. We find a strong, positive correlation where the outflow equivalent width (EWout) systematically increases with stellar mass at all cosmic epochs. This suggests strong absorption from extensive outflowing reservoirs in high-mass galaxies.
@@
    Our work provides the first statistical constraints on cool gas outflows in the low-mass ($M_{*}\lesssim 10^{9.5}~\mathrm{M}_{\odot}$), high-redshift ($z>3$) regime. These measurements provide a crucial baseline for feedback models in the numerous progenitors of typical present-day galaxies.
@@
    There is no significant evolution in the cool gas outflow properties at high redshifts ($z>3$). Both the scaling relation between outflow equivalent width and stellar mass, and the outflow velocities themselves, are remarkably consistent from $z\approx 3$ to $z\approx 10$. This suggests that a persistent, unevolving feedback mechanism is already governing the baryon cycle. This finding challenges theoretical models that predict a fundamental change or breakdown of feedback at Cosmic Dawn, such as the feedback-free starburst model, indicating it is likely not the dominant feedback process in the observed galaxy population.
Our results provide fundamental, statistically robust observational constraints on the interplay between galactic feedback and the baryon cycle over 12 billion years of cosmic history. These measurements of outflow scaling relations, particularly in the previously unexplored high-redshift ($z>3$), low-mass regime, will be crucial for calibrating the next generation of cosmological simulations and refining our theoretical understanding of galaxy evolution towards the Cosmic Dawn.
EW thanks support of the National Science Foundation of China (Nos. 12473008). CL is supported by the Fundamental Research Funds for the Central Universities (WK2030250123). The authors gratefully acknowledge the support of Cyrus Chun Ying Tang Foundations. The data products presented herein are retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.
@@ Appendix A Comparison with Equal-Weight Stacking
In addition to the S/N-weighted stacking method adopted in our main analysis (Section 3), we perform a comparative test using an unweighted (equal-weight) stacking approach. In this alternative scheme, every spectrum contributes equally to the final stack, regardless of its individual S/N.
The left panels of Figure 7 presents the results of this test. A direct visual comparison with our fiducial stacks (right panels, the same as the left panels of Figure 2) immediately reveals the inferiority of this simpler method. Through involving poorly normalized individual spectra, composite spectra derived from unweighted stacking exhibits larger noise which is non-physical. This may potentially wash out the subtle, underlying absorption features. As a result, the continuum is more poorly defined, and the overall statistical significance of the detection is substantially lower. This comparison therefore underscores the critical importance of our S/N-weighting scheme for robustly detecting and characterizing the faint outflow signals in our sample. Although the effect of weighting is more subtle in our current, relatively small JWST sample, its importance becomes dramatic when stacking larger datasets to recover very faint signals.
@@ Appendix B Individual Galaxy Spectra of Absorbers
1 cosmos-curti-v4_g140m-f100lp_1879_3747 $150.06155$ $2.23362$ 2.522 10.29 372.60-372.60 4.56
2 bluejay-north-v4_g140m-f100lp_1810_16474 $150.11344$ $2.34747$ 2.552 10.07 234.48-234.48 5.93
3 bluejay-south-v4_g140m-f100lp_1810_8888 $150.12873$ $2.26979$ 2.561 10.49 326.84-326.84 1.85
4 cosmos-curti-v4_g140m-f100lp_1879_1858 $150.08136$ $2.20749$ 2.658 9.65 178.03-178.03 -0.68
5 bluejay-south-v4_g140m-f100lp_1810_8002 $150.11993$ $2.26102$ 2.686 10.99 439.10-439.10 8.38
6 jades-gds03-v4_g140m-f070lp_1286_181082 $53.12218$ 27.85834-27.85834 2.805 10.29 206.80-206.80 9.54
7 jades-gds1-v4_g140m-f070lp_1286_42789 $53.18607$ 27.80225-27.80225 2.807 9.78 474.79-474.79 1.43
8 gds-udeep-v4_g140m-f070lp_3215_199773 $53.16324$ 27.80906-27.80906 2.821 10.72 481.77-481.77 3.81
9 bluejay-north-v4_g140m-f100lp_1810_18071 $150.13527$ $2.36361$ 2.822 10.81 1306.97-1306.97 0.69
10 jades-gdn2-v4_g140m-f070lp_1181_20185 $189.10652$ $62.23063$ 2.932 10.43 288.94-288.94 5.60
11 jades-gds04-v4_g140m-f070lp_1286_53237 $53.04188$ 27.86134-27.86134 2.971 10.02 273.17-273.17 3.86
12 jades-gdn10-v4_g140m-f070lp_1181_77071 $189.22890$ $62.19548$ 2.974 10.58 386.53-386.53 8.36
13 bluejay-north-v4_g140m-f100lp_1810_17793 $150.13779$ $2.36100$ 2.986 10.35 317.47-317.47 5.04
14 bluejay-north-v4_g140m-f100lp_1810_17906 $150.12949$ $2.36170$ 2.987 10.12 95.53-95.53 2.45
15 jades-gdn11-v4_g140m-f070lp_1181_25603 $189.17021$ $62.23287$ 2.989 10.26 8285.03-8285.03 0.06
16 aurora-gdn02-v4_g140m-f100lp_1914_22384 $189.16149$ $62.25202$ 2.993 10.29 311.04-311.04 3.62
17 aurora-gdn02-v4_g140m-f100lp_1914_19848 $189.16046$ $62.23943$ 2.993 10.01 355.90-355.90 3.07
18 jades-gds-wide-v4_g140m-f070lp_1180_11929 $53.13048$ 27.77596-27.77596 2.997 10.17 290.74-290.74 2.78
19 jades-gds-wide3-v4_g140m-f070lp_1180_197911 $53.16531$ 27.81414-27.81414 3.063 11.20 321.37-321.37 6.00
20 gds-rieke-v4_g140m-f100lp_1207_54612 $53.14468$ 27.77118-27.77118 3.083 9.68 $348.25$ -1.75
21 jades-gds1-v4_g140m-f070lp_1286_54612 $53.14468$ 27.77118-27.77118 3.083 9.68 131.75-131.75 -3.29
22 aurora-gdn01-v4_g140m-f100lp_1914_4113 $150.13793$ $2.22025$ 3.084 10.57 323.91-323.91 6.86
23 cosmos-curti-v4_g140m-f100lp_1879_599 $150.09334$ $2.18999$ 3.087 10.82 173.67-173.67 2.98
24 bluejay-south-v4_g140m-f100lp_1810_7934 $150.09731$ $2.26012$ 3.093 10.03 250.87-250.87 -2.07
25 aurora-gdn02-v4_g140m-f100lp_1914_21033 $189.21160$ $62.24571$ 3.111 9.37 410.13-410.13 2.04
26 jades-gdn11-v4_g140m-f070lp_1181_25351 $189.17337$ $62.23041$ 3.129 10.73 510.18-510.18 -0.35
27 aurora-gdn01-v4_g140m-f100lp_1914_4740 $150.15878$ $2.22606$ 3.157 10.51 $136.64$ 0.82
28 aurora-gdn01-v4_g140m-f100lp_1914_6143 $150.18320$ $2.24024$ 3.158 10.07 345.79-345.79 4.02
29 excels-uds04-v4_g140m-f100lp_3543_61762 $34.27260$ 5.22469-5.22469 3.227 9.98 298.31-298.31 1.75
30 aurora-gdn01-v4_g140m-f100lp_1914_8363 $150.13641$ $2.26468$ 3.247 9.83 775.53-775.53 -0.41
31 bluejay-south-v4_g140m-f100lp_1810_8338 $150.07943$ $2.26452$ 3.268 10.32 502.21-502.21 5.57
32 jades-gdn2-v4_g140m-f070lp_1181_2945 $189.14089$ $62.28066$ 3.336 10.34 465.04-465.04 3.27
33 aurora-gdn02-v4_g140m-f100lp_1914_11584 $189.10502$ $62.28033$ 3.362 11.04 233.73-233.73 4.09
34 aurora-gdn02-v4_g140m-f100lp_1914_23927 $189.24105$ $62.25951$ 3.363 9.15 139.23-139.23 -3.15
35 bluejay-south-v4_g140m-f100lp_1810_13297 $150.11868$ $2.31445$ 3.488 9.82 379.84-379.84 9.27
36 jades-gds03-v4_g140m-f070lp_1286_180657 $53.11252$ 27.85900-27.85900 3.583 9.18 $972.03$ 0.18
37 excels-uds04-v4_g140m-f100lp_3543_39063 $34.29044$ 5.26208-5.26208 3.700 10.71 261.85-261.85 1.85
38 excels-uds04-v4_g140m-f100lp_3543_56133 $34.30228$ 5.23394-5.23394 3.722 10.49 270.23-270.23 2.45
39 jades-gdn10-v4_g140m-f070lp_1181_82948 $189.23303$ $62.23678$ 3.907 10.82 206.98-206.98 -2.75
40 excels-uds02-v4_g140m-f100lp_3543_117004 $34.28827$ 5.13754-5.13754 4.081 10.62 276.72-276.72 5.53
41 excels-uds03-v4_g140m-f100lp_3543_45052 $34.36713$ 5.25202-5.25202 4.233 10.26 216.83-216.83 -4.80
42 aurora-gdn02-v4_g140m-f100lp_1914_100067 $189.23479$ $62.25748$ 5.188 9.12 422.16-422.16 1.08
@@
Note. — The outflow properties are measured using the boxcar method described in Section 3.
In typical outflows, $\rm EW_{out}>0$ or $v_{\rm out}<0$. Any result with $\rm EW_{out}<0$ or $v_{\rm out}>0$ is purely numerical outputs (see Section 3).
@@
These examples highlight the significant kinematic diversity present in the underlying population. The majority of the selected objects display clear blueshifted absorption profiles (negative velocities), providing direct visual confirmation of the outflow signatures that our stacking method robustly recovers. However, the gallery also reveals more complex cases. A few systems (e.g., No.34, No.39, marked by green color) show visually net redshifted absorption, a potential signature of cool gas accreting onto the galaxy (inflow). A comprehensive analysis of these individual systems, including detailed spectral profile fitting to disentangle the complex interplay of outflows, inflows, and virial motions, is beyond the scope of this statistically-focused Letter. We will present such an analysis in a forthcoming paper.

### v2

@@ IV.4 Outflow Velocities
V Discussion
V.1 The Fate of the Outflowing Gas
To assess the fate of the outflowing gas, we compare our measured outflow velocities ($v_{\rm out}$) with the escape velocities ($v_{\rm esc}$) of the host dark matter halos. We estimate the halo mass ($M_{h}$) for each galaxy using the redshift-dependent stellar-to-halo mass relation (SHMR) from Moster et al. (2013). The virial radius ($R_{\rm vir}$) is then calculated, where the virial overdensity $\Delta_{\rm vir}(z)$ is derived from the approximation of Bryan and Norman (1998). We define the circular velocity of the halo as $v_{\rm circ}\approx\sqrt{GM_{h}/R_{\rm vir}}$. Following previous studies (e.g., Veilleux et al., 2005, 2020; Weldon et al., 2024), we adopt a simplified approximation for the escape velocity: $v_{\rm esc}\approx 3\times v_{\rm circ}$. We caution that these $v_{\rm esc}$ curves are dependent on the assumed redshift evolution of the SHMR, here adopted from Moster et al. (2013). The SHMR at high redshifts ($z\gtrsim 4$) remains poorly constrained observationally and varies among different theoretical models and simulations (e.g., Behroozi et al., 2019; Girelli et al., 2020), introducing systematic uncertainties to the exact normalization of the escape velocity thresholds.
The resulting $M_{*}-v_{\rm esc}$ relations for different redshifts are plotted as colored solid curves in Figure 5. As shown in the figure, both our stacked measurements (star symbols) and the majority of individual detections (circles) lie below these escape velocity curves across the entire stellar mass range probed. This indicates that the cool gas outflows traced by Mg II are predominantly gravitationally bound to their host halos. Instead of escaping into the IGM, this material is likely to decelerate and eventually rain back onto the galaxy, consistent with a “galactic fountain” scenario where the baryon cycle is regulated by the recycling of chemically enriched gas. This result aligns with recent statistical studies of ionized gas kinematics using JWST (Cooper et al., 2025; Wu, 2025; Xu et al., 2025) that the average outflow velocities are generally insufficient to escape the galaxy potential.
Table 1: Properties of Stacked Samples and Outflow Measurements. The columns are as follows: (1) Redshift bin. (2) Weighted mean redshift of the bin. (3) Stellar mass bin. For rows with a dash, measurements correspond to the entire redshift bin. (4) Weighted mean of the logarithm of the stellar mass. (5) Number of spectra in the stack. (6) Weighted mean outflow velocity derived from the Mg II doublet. (7) Outflow equivalent width of the Mg II doublet. (8) Lower limit on mass outflow rate.
$z$ $\langle z\rangle_{w}$ $\log M_{*}$ $\langle\log M_{*}\rangle_{w}$ NN $v_{\rm out}$ $\rm{EW_{\rm out}}$ $\dot{M}_{\mathrm{out,min}}$
($\rm M_{\odot}$) ($\rm M_{\odot}$) ($\rm km\,s^{-1}$) (Å) ($\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$)
1–2 1.836 9.65 334 $481^{+395}_{-453}$ $54^{+0.70}_{-0.68} 0.0$890.089
1–2 1.828 8.0–9.5 8.96 224 $371^{+804}_{-591}$ $20^{+1.38}_{-1.23} 0.0$680.068
1–2 1.842 9.5–12.0 10.14 110 $382^{+105}_{-181}$ $83^{+0.77}_{-0.88} 0.0$700.070
2–3 2.602 9.47 763 $365^{+32}_{-30}$ $1.75^{+0.38}_{-0.36}$ $0.067$
2–3 2.608 8.0–9.5 8.99 527 $339^{+104}_{-112}$ $96^{+0.56}_{-0.49} 0.0$620.062
2–3 2.582 9.5–10.5 9.93 175 $364^{+33}_{-31}$ $2.21^{+0.42}_{-0.56}$ $0.067$
2–3 2.648 10.5–12.0 10.78 61 $350^{+33}_{-45}$ $3.92^{+1.38}_{-1.24}$ $0.065$
3–4 3.458 9.39 586 $260^{+23}_{-21}$ $2.02^{+0.32}_{-0.30}$ $0.048$
3–4 3.486 8.0–9.5 8.95 428 $265^{+53}_{-43}$ $1.44^{+0.43}_{-0.43}$ $0.049$
3–4 3.410 9.5–10.5 9.88 124 $249^{+40}_{-27}$ $1.89^{+0.54}_{-0.53}$ $0.046$
3–4 3.423 10.5–12.0 10.88 34 $314^{+38}_{-28}$ $4.75^{+0.96}_{-1.15}$ $0.058$
4–5 4.356 9.24 210 $400^{+84}_{-189}$ $1.26^{+0.61}_{-0.74}$ $0.074$
4–5 4.385 8.0–9.5 8.89 174 $358^{+108}_{-162}$ $1.23^{+0.69}_{-0.75}$ $0.066$
4–5 4.281 9.5–12.0 10.14 36 $381^{+241}_{-256}$ $1.30^{+1.32}_{-1.44}$ $0.070$
5–10 6.067 9.01 497 $336^{+57}_{-85}$ $1.81^{+0.75}_{-0.70}$ $0.062$
5–10 6.078 8.0–9.5 8.81 450 $323^{+98}_{-98}$ $1.23^{+0.76}_{-0.71}$ $0.060$
5–10 6.017 9.5–12.0 9.84 47 $336^{+66}_{-91}$ $3.18^{+1.62}_{-1.85}$ $0.062$
[…]
V.2 Lower Limits on Mass Outflow Rates
[…]
Based on the conservative assumptions of a minimal launch radius ($r_{\rm min}=1$ kpc) and a saturated column density limit ($N_{\rm H,min}\approx 2.5\times 10^{18}\,{\rm cm}^{-2}$), the mass outflow rate lower limit is estimated by scaling the standard relation (Weiner et al., 2009):
$\dot{M}_{\rm out,min}=22\,{\rm M}_{\odot}\,{\rm yr}^{-1}\cdot C_{f}\cdot\frac{N_{\rm H,min}}{10^{20}\,{\rm cm}^{-2}}\cdot\frac{1\,{\rm kpc}}{5\,{\rm kpc}}\cdot\frac{v_{\rm out}}{300\,{\rm km\,s}^{-1}},$ (7)
where $C_{f}\approx 0.5$(Pessa et al., 2024) is the covering fraction and $N_{\rm H,min}$ is derived from $N({\rm Mg\,II})>10^{14}\,{\rm cm}^{-2}$ assuming solar abundance. These derived lower limits for stackted results are listed in Table 1. We note that these values are strictly minimal constraints; realistic mass outflow rates, accounting for larger radii ($R\gtrsim 5$ kpc) and ionization corrections (Mg II fraction $f\lesssim 0.1$), are likely an order of magnitude higher.
V.3 Synergy of Stacking and Individual Analysis
@@
A key strength of our work is the synergistic combination of a statistical stacking analysis with a targeted study of individual detections. As illustrated in Figure 7 (Appendix B) , the galaxies with the most prominent, high-S/N absorption features are typically massive ($M_{*}>10^{10}~\rm{M_{\odot}}$). These individual detections serve as unambiguous proof-of-concept, confirming the physical reality of the outflow phenomenon and revealing the rich diversity of absorption profile morphologies. This diversity includes not only powerful outflows but also a few intriguing cases of redshifted absorption as potential signatures of gas accretion (inflow), which we defer to a future, more detailed analysis (see Appendix B).
However, these massive objects represent only a small fraction of the galaxy population. By extending our analysis to lower stellar masses ($M_{*}<10^{9.5}M_{\odot}$) via spectral stacking, we demonstrate that cool gas outflows are not limited to rare, massive starbursts but are a ubiquitous feature of typical galaxies at these epochs. The combination of individual characterization and statistical stacking thus provides a complete picture of feedback across the full dynamic range of our sample.
VI Summary
@@
We present the first statistical study of Mg II-traced cool gas outflows in galaxies spanning a wide cosmic timeline from $z\approx 1$ to $z>6$ (with sample coverage extending to $z\sim 10$). Leveraging thousands of public JWST/NIRSpec spectra, we employ a S/N-weighted spectral stacking technique, complemented by an analysis of individual high-S/N detections, to characterize the faint absorption signatures of the Mg II doublet. Our sample is shown to be broadly representative of the underlying galaxy population. Our main findings are summarized as follows:
@@
    We robustly detect blueshifted Mg II absorption, indicative of cool galactic outflows, in nearly all redshift and stellar mass bins from $z\approx 1$ to $z>6$, with the exception of the lowest-mass bin ($\log M_{*}/M_{\odot}<9.5$) at $z<2$ where the signal is marginal. The derived typical outflow velocities ($\sim 300-400$ km s-1) suggest that cool gas outflows are a common feature of galaxies across the probed cosmic history.
@@
    The properties of the cool gas outflows appear correlated with the stellar mass of the host galaxy. We observe that the outflow equivalent width (EWout) systematically increases with stellar mass at all cosmic epochs (with the fitted slope of $1.21\pm 0.35$). This trend suggests that more massive galaxies are associated with stronger absorption from cool, outflowing gas reservoirs.
@@
    We find no evidence for significant evolution in the outflow scaling relations at high redshifts ($z>3$). Both the relation between EWout and stellar mass, as well as the outflow velocities, are broadly consistent from $z\approx 3$ to $z\approx 6$ (where our highest redshift bin is centered). These consistencies suggest that the feedback mechanisms governing the cool gas phase operate efficiently even at the epoch of cosmic dawn.
@@
    The average outflow velocities measured in our stacks generally remain below the estimated escape velocities of the host dark matter halos. This indicates that the bulk of the cool gas traced by Mg II is likely gravitationally bound and may eventually decelerate and return to the galaxy. These results are consistent with a “galactic fountain” scenario, where the baryon cycle is regulated by the recycling of chemically enriched gas rather than by efficient removal into the intergalactic medium.
Our results provide fundamental, statistically observational constraints on the interplay between galactic feedback and the baryon cycle over 12 billion years of cosmic history. By characterizing outflow scaling relations in the high-redshift ($z>3$), low-mass regime, we provide statistical constraints on the cool gas phase that were previously lacking, complementing recent studies of ionized outflows (e.g., Carniani et al., 2024; Cooper et al., 2025; Xu et al., 2025). These findings will be crucial for calibrating the next generation of cosmological simulations and refining our theoretical understanding of galaxy evolution towards cosmic dawn.
The authors thank the anonymous referee for the helpful comments that improved the quality of this Letter. EW thanks support of the National Science Foundation of China (Nos. 12473008). CL is supported by the Fundamental Research Funds for the Central Universities (WK2030250123). The authors gratefully acknowledge the support of Cyrus Chun Ying Tang Foundations. The data products presented herein are retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.
@@ Appendix A Comparison with Equal-Weight Stacking
In addition to the S/N-weighted stacking method adopted in our main analysis (Section III), we perform a comparative test using an unweighted (equal-weight) stacking approach. In this alternative scheme, every spectrum contributes equally to the final stack, regardless of its individual S/N.
The left panels of Figure 6 presents the results of this test. A direct visual comparison with our fiducial stacks (right panels, the same as the left panels of Figure 2) immediately reveals the inferiority of this simpler method. Through involving poorly normalized individual spectra, composite spectra derived from unweighted stacking exhibit larger noise which is non-physical. This may potentially wash out the subtle, underlying absorption features. As a result, the continuum is more poorly defined, and the overall statistical significance of the detection is substantially lower. This comparison therefore underscores the critical importance of our S/N-weighting scheme for robustly detecting and characterizing the faint outflow signals in our sample. Although the effect of weighting is more subtle in our current, relatively small JWST sample, its importance becomes dramatic when stacking larger datasets to recover very faint signals.
@@ Appendix B Individual Galaxy Spectra of Absorbers
1 cosmos-curti-v4_g140m-f100lp_1879_3747 $150.06155$ $2.23362$ 2.522 10.29 $372^{+40}_{-39}$ $4.51^{+1.02}_{-1.02}$
2 bluejay-north-v4_g140m-f100lp_1810_16474 $150.11344$ $2.34747$ 2.552 10.07 $234^{+80}_{-58}$ $5.99^{+1.82}_{-1.75}$
3 bluejay-south-v4_g140m-f100lp_1810_8888 $150.12873$ $2.26979$ 2.561 10.49 $347^{+199}_{-179}$ $1.86^{+1.92}_{-2.01}$
4 cosmos-curti-v4_g140m-f100lp_1879_1858 $150.08136$ $2.20749$ 2.658 9.65 $302^{+276}_{-311}$ $0.72^{+0.94}_{-0.94}$
5 bluejay-south-v4_g140m-f100lp_1810_8002 $150.11993$ $2.26102$ 2.686 10.99 $439^{+21}_{-22}$ $8.41^{+1.21}_{-1.13}$
6 jades-gds03-v4_g140m-f070lp_1286_181082 $53.12218$ 27.85834-27.85834 2.805 10.29 $208^{+58}_{-42}$ $9.57^{+2.45}_{-2.30}$
7 jades-gds1-v4_g140m-f070lp_1286_42789 $53.18607$ 27.80225-27.80225 2.807 9.78 $394^{+247}_{-358}$ $1.42^{+1.95}_{-1.72}$
8 gds-udeep-v4_g140m-f070lp_3215_199773 $53.16324$ 27.80906-27.80906 2.821 10.72 $468^{+67}_{-144}$ $3.84^{+1.80}_{-1.96}$
9 bluejay-north-v4_g140m-f100lp_1810_18071 $150.13527$ $2.36361$ 2.822 10.81 $737^{+1296}_{-955}$ $0.63^{+1.24}_{-1.19}$
10 jades-gdn2-v4_g140m-f070lp_1181_20185 $189.10652$ $62.23063$ 2.932 10.43 $293^{+115}_{-117}$ $5.58^{+2.90}_{-2.98}$
11 jades-gds04-v4_g140m-f070lp_1286_53237 $53.04188$ 27.86134-27.86134 2.971 10.02 $280^{+152}_{-86}$ $3.91^{+1.91}_{-1.87}$
12 jades-gdn10-v4_g140m-f070lp_1181_77071 $189.22890$ $62.19548$ 2.974 10.58 $386^{+39}_{-53}$ $8.32^{+2.36}_{-2.26}$
13 bluejay-north-v4_g140m-f100lp_1810_17793 $150.13779$ $2.36100$ 2.986 10.35 $314^{+40}_{-36}$ $5.10^{+0.95}_{-1.12}$
14 bluejay-north-v4_g140m-f100lp_1810_17906 $150.12949$ $2.36170$ 2.987 10.12 $91^{+159}_{-88}$ $2.42^{+1.06}_{-0.99}$
15 jades-gdn11-v4_g140m-f070lp_1181_25603 $189.17021$ $62.23287$ 2.989 10.26 $514^{+1683}_{-1014}$ $0.04^{+0.87}_{-0.90}$
16 aurora-gdn02-v4_g140m-f100lp_1914_22384 $189.16149$ $62.25202$ 2.993 10.29 $314^{+56}_{-46}$ $3.59^{+0.96}_{-0.89}$
17 aurora-gdn02-v4_g140m-f100lp_1914_19848 $189.16046$ $62.23943$ 2.993 10.01 $356^{+18}_{-21}$ $3.08^{+0.39}_{-0.41}$
18 jades-gds-wide-v4_g140m-f070lp_1180_11929 $53.13048$ 27.77596-27.77596 2.997 10.17 $310^{+182}_{-153}$ $2.67^{+2.16}_{-2.06}$
19 jades-gds-wide3-v4_g140m-f070lp_1180_197911 $53.16531$ 27.81414-27.81414 3.063 11.20 $321^{+32}_{-30}$ $5.94^{+1.13}_{-1.13}$
20 gds-rieke-v4_g140m-f100lp_1207_54612 $53.14468$ 27.77118-27.77118 3.083 9.68 $170^{+824}_{-313}$ $1.74^{+1.52}_{-1.47}$
21 jades-gds1-v4_g140m-f070lp_1286_54612 $53.14468$ 27.77118-27.77118 3.083 9.68 $135^{+176}_{-92}$ $3.24^{+1.27}_{-1.36}$
22 aurora-gdn01-v4_g140m-f100lp_1914_4113 $150.13793$ $2.22025$ 3.084 10.57 $322^{+21}_{-20}$ $6.86^{+0.92}_{-0.91}$
23 cosmos-curti-v4_g140m-f100lp_1879_599 $150.09334$ $2.18999$ 3.087 10.82 $173^{+141}_{-78}$ $2.97^{+1.22}_{-1.11}$
24 bluejay-south-v4_g140m-f100lp_1810_7934 $150.09731$ $2.26012$ 3.093 10.03 $293^{+249}_{-186}$ $1.97^{+2.18}_{-2.48}$
25 aurora-gdn02-v4_g140m-f100lp_1914_21033 $189.21160$ $62.24571$ 3.111 9.37 $410^{+65}_{-62}$ $2.06^{+0.61}_{-0.67}$
26 jades-gdn11-v4_g140m-f070lp_1181_25351 $189.17337$ $62.23041$ 3.129 10.73 $285^{+336}_{-280}$ $0.38^{+2.61}_{-2.64}$
27 aurora-gdn01-v4_g140m-f100lp_1914_4740 $150.15878$ $2.22606$ 3.157 10.51 $59^{+524}_{-206}$ $0.81^{+0.58}_{-0.62}$
28 aurora-gdn01-v4_g140m-f100lp_1914_6143 $150.18320$ $2.24024$ 3.158 10.07 $348^{+30}_{-28}$ $4.02^{+0.65}_{-0.72}$
29 excels-uds04-v4_g140m-f100lp_3543_61762 $34.27260$ 5.22469-5.22469 3.227 9.98 $322^{+220}_{-244}$ $1.80^{+2.02}_{-2.05}$
30 aurora-gdn01-v4_g140m-f100lp_1914_8363 $150.13641$ $2.26468$ 3.247 9.83 $420^{+623}_{-359}$ $0.38^{+0.82}_{-0.79}$
31 bluejay-south-v4_g140m-f100lp_1810_8338 $150.07943$ $2.26452$ 3.268 10.32 $505^{+62}_{-82}$ $5.48^{+1.62}_{-1.43}$
32 jades-gdn2-v4_g140m-f070lp_1181_2945 $189.14089$ $62.28066$ 3.336 10.34 $451^{+100}_{-176}$ $3.36^{+1.99}_{-1.94}$
33 aurora-gdn02-v4_g140m-f100lp_1914_11584 $189.10502$ $62.28033$ 3.362 11.04 $232^{+47}_{-40}$ $4.10^{+0.89}_{-0.90}$
34 aurora-gdn02-v4_g140m-f100lp_1914_23927 $189.24105$ $62.25951$ 3.363 9.15 $157^{+249}_{-121}$ $3.14^{+1.92}_{-2.08}$
35 bluejay-south-v4_g140m-f100lp_1810_13297 $150.11868$ $2.31445$ 3.488 9.82 $379^{+30}_{-34}$ $9.24^{+1.68}_{-1.68}$
36 jades-gds03-v4_g140m-f070lp_1286_180657 $53.11252$ 27.85900-27.85900 3.583 9.18 $347^{+327}_{-357}$ $0.17^{+2.41}_{-2.54}$
37 excels-uds04-v4_g140m-f100lp_3543_39063 $34.29044$ 5.26208-5.26208 3.700 10.71 $307^{+306}_{-292}$ $1.82^{+3.04}_{-2.94}$
38 excels-uds04-v4_g140m-f100lp_3543_56133 $34.30228$ 5.23394-5.23394 3.722 10.49 $272^{+108}_{-78}$ $2.38^{+1.26}_{-1.22}$
39 jades-gdn10-v4_g140m-f070lp_1181_82948 $189.23303$ $62.23678$ 3.907 10.82 $210^{+80}_{-50}$ $2.73^{+0.66}_{-0.73}$
40 excels-uds02-v4_g140m-f100lp_3543_117004 $34.28827$ 5.13754-5.13754 4.081 10.62 $277^{+35}_{-37}$ $5.52^{+1.18}_{-1.23}$
41 excels-uds03-v4_g140m-f100lp_3543_45052 $34.36713$ 5.25202-5.25202 4.233 10.26 $216^{+138}_{-83}$ $4.76^{+2.13}_{-2.04}$
42 aurora-gdn02-v4_g140m-f100lp_1914_100067 $189.23479$ $62.25748$ 5.188 9.12 $396^{+265}_{-297}$ $1.09^{+1.79}_{-1.65}$
@@
Note. — The outflow properties are measured using the boxcar method described in Section III.
Typical outflows correspond to $\rm EW_{out}>0$ and $v_{\rm out}>0$. Any result with $\rm EW_{out}<0$ or $v_{\rm out}<0$ is purely numerical outputs (see Section III). Note that these negative values do not necessarily indicate gas inflows, as our measurement methodology is not designed to characterize inflow kinematics.
@@
These examples highlight the significant kinematic diversity present in the underlying population. The majority of the selected objects display clear blueshifted absorption profiles, providing direct visual confirmation of the outflow signatures that our stacking method robustly recovers. However, the gallery also reveals more complex cases. A few systems (e.g., No.34, No.39, marked by green color) show visually net redshifted absorption, a potential signature of cool gas accreting onto the galaxy (inflow). A comprehensive analysis of these individual systems, including detailed spectral profile fitting to disentangle the complex interplay of outflows, inflows, and virial motions, is beyond the scope of this statistically-focused Letter. We will present such an analysis in a forthcoming paper.
@@
Appendix C The Cosmic Evolution of Outflow Velocity
Figure 8: The cosmic evolution of outflow velocity. Our measurements from the full redshift bins stacks are shown as large star symbols. The color of all data points corresponds to the weighted mean stellar mass of the respective galaxy sample, as indicated by the color bar on the right. For context, we compare our results with a compilation of literature data from Weiner et al. (2009), Bordoloi et al. (2014),Sugahara et al. (2019), Davis et al. (2023), Yu et al. (2025), Liboni et al. (2025), Valentino et al. (2025),Cooper et al. (2025),Wu (2025) and Xu et al. (2025).
Figure 8 presents another view of Figure 5, showing the outflow velocity as a function of cosmic time. At first glance, the evolution of our measured outflow velocities (star symbols) with redshift appears complex, with no simple monotonic trend. The resulting best-fit relation is marked in the legend. However, this is primarily a consequence of selection effects, as the mean stellar mass of our stacked samples generally decreases with increasing redshift, which is clearly visualized by the color-coding of the data points. Nevertheless, this analysis highlights the unique contribution of our work: providing statistically derived outflow velocities for typical star-forming galaxies at $z>3$ and approaching cosmic dawn.